# autocorrelation function CFR channel frequency response

### v3

@@ III-B Relation Among L, M, and N
The maximum length $L$, the maximum number of paths $N$, and the number of frequency samples $M$ depend among each other and must satisfy certain constraints. Since only a discrete and finite number of frequencies is available, the expression in (1) given for the continuous frequency domain can be rewritten for the discrete domain as
@@
$\Delta f=\frac{f_{M-1}-f_{0}}{M-1}=\frac{1}{N\frac{L}{N\nu}}=\frac{\nu}{L}\Rightarrow L=\frac{(M-1)\nu}{f_{M-1}-f_{0}}$ (12)
Since $f_{M-1}=79.9358$ MHz and $f_{0}=1$ MHz, $M=1262$ and the propagation speed is $\nu=2\cdot 10^{8}$m/s, the maximum possible length that can be assumed is $L=3195$ m. The limit of more than 3 kilometers turns out to be a maximum length that is long enough to describe any generic in-home network topology. Moreover, always referring to the DFT representation of the CFR reported in (11), the DFT frequency period, i.e., $N\Delta f$, must be greater or equal than two times the maximum channel frequency, i.e., $f_{M-1}$, in order to avoid aliasing problems according to the sampling theorem. Thus, the relationship among $N$ and $L$ turns out to be
$N\Delta f=\frac{1}{\tau}\geq 2f_{M-1}\Rightarrow N\geq\frac{2f_{M-1}L}{\nu}.$ (13)
If the aforesaid values for $f_{M-1}$ and $\nu$ are considered, the relation in (13) becomes $N\geq 0.8L$. A final relation among $N$ and $L$ can be noticed by rewriting the imaginary part of the exponential term in (1) as
@@
$d_{1}=\frac{L}{N}\geq\frac{\lambda_{min}}{2}\Rightarrow N\leq\frac{2f_{M-1}L}{\nu}$ (15)
It can be noted that, given the above mentioned specifications, as well as the relations (13) and (15), the relationship among $L$ and $N$ must hold with equality, i.e., $N=0.8L=2554$ for the considered database.
IV Analysis of the MPM Parameters
This section analyzes the statistical properties of the MPM parameters obtained according to the fitting procedure detailed in Section III for the overall set of measured channels. First, the relation of some model parameters to the main magnitudes used to characterize the channel response, such as the delay spread and the average channel gain. Then, the PDF of the MPM parameters is estimated and, finally, the statistical relation between the model parameters is studied.
@@ IV-A Relation to the Channel Characteristics
The parameters $a_{0}$ and $a_{1}$ where obtained from the robust regression fit of the amplitude response of the measured channels, with $a_{1}$ being related to the slope of the attenuation profile and $a_{0}$ with the y-intercept. Hence, the latter parameter has to be somehow related to the average channel gain, defined as $G\textrm{ (dB)}=\frac{1}{M}\sum_{m=0}^{M-1}20\log_{10}{|H(f_{m})|}$. This relation can be clearly observed in Fig. 2 (a), where the scatter plot of $a_{0}$ vs $G$ (dB) is depicted, along with its robust fit regression line, whose expression is given in Table II. As expected, channels with larger average gains need lower values of $a_{0}$ to be synthesized.
Figure 2: Scatter plot of (a) the value of $\bm{a_{0}}$ vs the average channel gain of each measured channel, $\bm{G}$ (dB); (b) the number of dominant paths vs the average gain and the delay spread of the channels; (c) value of $\bm{A}$ vs the average channel gain of each measured channel.
The number of dominant paths needed to fit a given CFR is related to the frequency selectivity of the channel, e.g., a single path suffices to synthesize a flat channel. The frequency selectivity of a channel is usually quantified by means of the coherence bandwidth or its inversely related magnitude, the delay spread (also referred to as multipath spread) [31]. In PLC these magnitudes are known to be related to the average channel gain [8]: channels with lower average channel gains tend to have larger delay spreads. Hence, these magnitudes are expected to be related to the number of dominant paths, $N$. Fig. 2 (b) illustrates this relation by showing the scatter plot of the number of dominant paths vs the average channel gain and the delay spread of the corresponding measured channel. Its robust regression plane, whose expression is given in Table II, is also depicted for reference. The goodness of this fit is confirmed by the NRMSE between the actual and predicted values of $N$, which is $-22$ dB.
The value of the normalization parameter $A$ is also related to the average channel gain, as shown in Fig. 2 (c), which depicts the scatter plot of $A$ vs $G$ (dB) and its corresponding exponential fitting function. The parameters of the latter, which have been obtained by means of a LS fitting, are given in Table II. As observed, the lower the average channel gain, the lower the value of $A$ needed to synthesize the channel.
Table II: Parameters of the fitting expressions displayed in Fig. 2.
@@
[…]
$\widehat{A}=\alpha\cdot e^{\beta\cdot G\textrm{ (dB)}}$ $8.2517\cdot 10^{-1}$ $1.0636\cdot 10^{-1}$
@@ IV-B Statistical Characterization
This subsection estimates the chief moments and the PDF of the MPM parameters, computed for the overall set of measured channels, that have been obtained using the fitting procedure presented in Section IV. For each empirical (estimated) PDF, the probability distribution that gives the best fit is given. To this end, the most common distribution functions are considered: beta, Birnbaum-Saunders, exponential, gamma, generalized extreme value (GEV), Gumbel, inverse Gaussian, logistic, log-logistic, log-normal, Nakagami, normal, Poisson, Rayleigh, Rician, $t$ location-scale and Weibull. The parameters of each distribution that yield the best fit to the set of MPM parameters are obtained by means of the maximum likelihood (ML) estimation assuming independent samples [32]. Afterwards, the Anderson-Darling goodness-of-fit test is applied and the distribution yielding the highest $p$-value is selected. When the test gives the same $p$-value for all distributions or the null hypothesis is rejected for all of them, the one with the highest value of the likelihood function is selected. For some parameters, more elaborate models consisting of mixture distributions have been considered.
The mean and standard deviation of the MPM parameters are given in Table III. The fitting statistical distributions with their corresponding parameter values are summarized in Table IV. These reported results allow future replicability of this study and simplify the development of improved statistical models.
Table III: Mean and standard deviation of the MPM parameters
@@
Figure 3: Empirical and fitted PDFs of the attenuation coefficients $\bm{a_{0}}$ and $\bm{a_{1}}$ and the normalization coefficient $\bm{A}$.
@@
Figure 4: Empirical and fitted PDFs of the number of dominant paths, $\bm{N}$, the modulus of the path gains $\bm{|g_{i}|}$ and the path lengths $\bm{d_{i}}$.
@@
The empirical s of the MPM parameters are reported in Fig. 3 and Fig. 4. Regarding $a_{0}$ and $a_{1}$, whose EPDFs are show in Fig. 3 (a) and (b), respectively, the distribution that gives the best fit is the GEV ($p$-value$=0.286$ for $a_{0}$ and 0.8941 for $a_{1}$), although the normal distribution also gives a high $p$-value (0.1967 for $a_{0}$ and 0.5664 for $a_{1}$). In practice, the GEV is a family of three types of extreme value distributions, namely type I, or Gumbel (or log-Weibull) [33], type II, or Fréchet, and type III, or Weibull distribution. This result makes sense since the GEV is the limit distribution of properly normalized maxima of a sequence of independent and identically distributed random variables. The GEV distribution for $a_{0}$ is justified since it represents the maximum (or the minimum in case of negative values of $a_{1}$) of the channel attenuation profile for the identified trend of each measured CFR.
@@
Regarding the normalization coefficient, $A$, whose EPDF is displayed in Fig. 3 (c), the Log-logistic gives the largest $p$-value (0.909) although, although the GEV and the log-normal also give high values ($p$-values $0.788$ and $0.4536$, respectively). The latter is not surprising, as $A$ is determined as the maximum value of the path gains, $|g_{i}|$, of each fitted channel.
@@
Regarding the probability distribution of the number of dominant paths, $N$, while this magnitude can only take non-negative integer values, discrete distributions like the binomial and Poisson ones give very poor fitting to the actual set of values. Hence, $N$ has been considered as a discretized version of a continuous random variable. This same approach has been also applied to fit the path length, $d_{i}$. By doing so, the GEV gives the best fit ($p$-value$=0.86752$) to the number of dominant paths, $N$, whose EPDF is displayed in Fig. 4 (a). As a matter of fact, $N$ represents the maximum number of dominant paths that actively contribute to reconstruct each measured CFR.
@@
Concerning the characterization of the path gain, its absolute value, $|g_{i}|$, is considered for the sake of simplicity, as the EPDF of $g_{i}$ shows a symmetric behavior around the $y$-axis. The statistical characterization of $|g_{i}|$ must take into account that the normalization process used to force $g_{i}\in[-1,1]$ in each fitted channel causes the probability of $|g_{i}|=1$ to be non-zero. Hence, $|g_{i}|$ has to be modeled by a mixed random variable (RV) consisting of a continuous component and a discrete term at $|g_{i}|=1$. Fig. 4 (b) shows that the continuous term of the EPDF can be fitted by a log-normal RV. Accordingly, the overall PDF can be expressed as $f(|g_{i}|)=\pi_{0}q(|g_{i}|\big{|}\mu,\sigma)+\pi_{1}\delta(|g_{i}|-1)$, where $q(\cdot\big{|}\mu,\sigma)$ denotes the PDF of a log-normal RV with parameters $\mu$ and $\sigma$ and $\delta(\cdot)$ is the Dirac delta function. The value of these parameters are given in Table IV.
@@
The EPDF of $d_{i}$, displayed in Fig. 4 (c), exhibits a bimodal behavior, with maxima around $d_{i}=50$ m and $d_{i}=3195$ m. The latter is an effect of the use of the discrete frequency domain version of (1), which obliges to set a limit for the largest path length. Since the observed bimodal behavior cannot be obtained with any of the distributions indicated above, we have resorted to mixture models. It has been found that the Weibull distribution gives a good fitting to the values of $d_{i}\leq 1500$ while the GEV, shifted and reflected about $d_{N-1}=\frac{N-1}{N}L$, adequately fits the values of $d_{i}>1500$. Hence, the PDF of $d_{i}$ can be modeled as $f(d_{i})=\pi_{0}g_{0}(d_{i}\big{|}\lambda_{0},k_{0})+\pi_{1}g_{1}(d_{N-1}-d_{i}\big{|}k_{1},\sigma_{1},\mu_{1})$, where $g_{0}(\cdot\big{|}\lambda_{0},k_{0})$ denotes the PDF of a Weibull distribution with scale and shape parameters $\lambda_{0}$ and $k_{0}$, respectively, and $g_{1}(\cdot\big{|}k_{1},\sigma_{1},\mu_{1})$ is the PDF of a GEV distribution with location, scale and shape parameters $k_{1}$, $\sigma_{1}$ and $\mu_{1}$, respectively, whose values are given in Table IV. As shown in Fig. 4 (c), the PDF of the proposed mixture gives a good fitting to the empirical one.
Table IV: Fitting distributions of the MPM parameters
@@
In order to get an in-depth analysis of the behavior of $d_{i}$, Fig. 5 (a) displays the path length, $d_{i}$, vs the dominant path index, $i$, for some illustrative measured channels. For the sake of clarity, the markers corresponding the same channel have been linked by a dashed line of the the same color. As seen, all channels have a set of paths whose length is uniformly distributed from zero (or almost zero) and a given value that depends on the channel. However, some channels, like Ch. 3-7, also have a set of paths with very large length. In many cases, like Ch. 3 and Ch. 4, there are no paths with intermediate distances. This behavior is related to the frequency selectivity of the channel: the larger the frequency selectivity, the large the distances of the paths required to synthesize it. This can be observed in Fig. 5 (b), where the scatter plot of the median value of the dominant path lengths vs the delay spread of each measured channel is given. As observed, channels with higher delay spread values (larger frequency selectivity) need larger paths to be synthesized. Futhermore, it is also expected that the higher the number of dominant paths, the larger the median length of the paths. Fig. 5 (c) confirms this end by showing a strong linear relationship between them.

### v4

@@ III-B Relation Among L, M, and N
The maximum length $L$, the maximum number of paths $N$ depend on the considered frequency band and the frequency resolution of the measured CFR. Hence, they are also related to the number of frequency samples, $M$, of the latter and must satisfy certain constraints which are now derived.
Since only a discrete and finite number of frequencies is available, the expression in (1) given for the continuous frequency domain can be rewritten for the discrete domain as
@@
$\Delta f=\frac{f_{M-1}-f_{0}}{M-1}=\frac{1}{N\frac{L}{N\nu}}\Rightarrow L=\frac{\nu}{\frac{f_{M-1}-f_{0}}{M-1}}=\frac{\nu}{\Delta f}.$ (12)
Expression (12) can be also derived from physical considerations. The time it takes for an input impulse to the channel to reach the receiver following the longest path (i.e., the one with the largest number of reflections) is $\tau=L/\nu$. The maximum duration of the channel impulse response that can be obtained from a sampled version of the CFR at multiples of $\Delta f$ while avoiding time-domain aliasing is $\tau=1/\Delta f$. Equating both expressions yields $L=\nu/\Delta f$.
For $f_{M-1}=79.9358$ MHz and $f_{0}=1$ MHz, $M=1262$ and the propagation speed is $\nu=2\cdot 10^{8}$m/s, the maximum possible length that can be assumed is $L=3195$ m. Since physical distances in indoor PLC networks are on the order of tens or few hundreds of meters, the limit of over $3$ kilometers is large enough to account for the longest distance traveled by all forward- and backward-propagating waves received with significant power.
Referring to the DFT representation of the CFR reported in (11), the DFT frequency period, i.e., $N\Delta f$, must be greater or equal than two times the maximum channel frequency, i.e., $f_{M-1}$, in order to avoid aliasing. Thus, the relationship among $N$ and $L$ turns out to be
$N\Delta f=\frac{1}{\tau}\geq 2f_{M-1}\Rightarrow N\geq\frac{2f_{M-1}L}{\nu}=\frac{f_{M-1}}{\Delta f}.$ (13)
As seen, $N$ is determined by the considered bandwidth and the frequency resolution of the measurements. If the aforesaid values for $f_{M-1}$ and $\nu$ are considered, the relation in (13) becomes $N\geq 0.8L$. A final relation among $N$ and $L$ can be noticed by rewriting the imaginary part of the exponential term in (1) as
@@
$d_{1}=\frac{L}{N}\geq\frac{\lambda_{min}}{2}\Rightarrow N\leq\frac{2f_{M-1}L}{\nu}=\frac{f_{M-1}}{\Delta f}.$ (15)
Expression (15) can be also derived from physical considerations. Assuming uniform distances, the time difference between the closest received paths is $\frac{L/N}{\nu}$. In order for these paths to be resolvable, this time must be greater or equal than the minimum sampling period, $T_{s}=\frac{1}{2f_{M-1}}$. This yields $\frac{L/N}{\nu}\geq\frac{1}{2f_{M-1}}$, which equals (15).
Given the specifications of the measured CFR, as well as the relations (13) and (15), the relationship among $L$ and $N$ must hold with equality, i.e., $N=0.8L=2554$ for the considered database.
IV Statistical Analysis of the MPM Parameters
This section analyzes the statistical properties of the MPM parameters obtained according to the fitting procedure detailed in Section III for the overall set of measured channels. First, the relation of some model parameters to the main quantities used to characterize the channel response, such as the delay spread and the average channel gain. Then, the PDF of the MPM parameters is estimated and, finally, the statistical relation between the model parameters is studied.
@@ IV-A Relation to the Channel Characteristics
The parameters $a_{0}$ and $a_{1}$ were obtained from the regression fit of the amplitude response of the measured channels, with $a_{1}$ being related to the slope of the attenuation profile and $a_{0}$ with the y-intercept. Hence, the latter parameter has to be somehow related to the average channel gain, defined as $G\textrm{ (dB)}=\frac{1}{M}\sum_{m=0}^{M-1}20\log_{10}{|H(f_{m})|}$}. This relation can be clearly observed in Fig. 3 (a), where the scatter plot of $a_{0}$ vs $G\;\textrm{(dB)}$} is depicted, along with its robust fit regression line, whose expression is given in Table III. As expected, channels with larger average gains need lower values of $a_{0}$ to be synthesized.
Figure 3: Scatter plot of (a) the value of $a_{0}$ vs the average channel gain of each measured channel, $G$ (dB); (b) the number of dominant paths vs the average gain and the delay spread of the channels; (c) value of $A$ vs the average channel gain of each measured channel.
The number of dominant paths needed to fit a given CFR is related to the frequency selectivity of the channel, e.g., a single path suffices to synthesize a flat channel. The frequency selectivity of a channel is usually quantified by means of the coherence bandwidth or its inversely related magnitude, the delay spread (also referred to as multipath spread) [Molish2011]. In PLC these magnitudes are known to be related to the average channel gain [Cortes23bis]: channels with lower average channel gains tend to have larger delay spreads. Hence, these magnitudes are expected to be related to the number of dominant paths, $N$. Fig. 3 (b) illustrates this relation by showing the scatter plot of the number of dominant paths vs the average channel gain and the delay spread of the corresponding measured channel. Its robust regression plane, whose expression is given in Table III, is also depicted for reference. The goodness of this fit is confirmed by the NRMSE between the actual and predicted values of $N$, which is 22-22 dB.
The value of the normalization parameter $A$ is also related to the average channel gain, as shown in Fig. 3 (c), which depicts the scatter plot of $A$ vs $G$ (dB) and its corresponding exponential fitting function. The parameters of the latter, which have been obtained by means of a LS fitting, are given in Table III. As observed, the lower the average channel gain, the lower the value of $A$ needed to synthesize the channel.
Table III: Parameters of the fitting expressions displayed in Fig. 3.
@@
$\widehat{a}_{0}=\alpha+\beta\cdot G\textrm{ (dB)}$ $1.8669\cdot 10^{-4}$ $3.4066\cdot 10^{-5} -$
$\widehat{N}=\alpha+\beta\cdot ds\textrm{ ($\mu$s)}+\gamma\cdot G\textrm{ (dB)}$ $40.4009$ $185.7535$ $4.5097$
$\widehat{A}=\alpha\cdot e^{\beta\cdot G\textrm{ (dB)}}$ $8.2517\cdot 10^{-1}$ $1.0636\cdot 10^{-1}$ -
@@ IV-B Statistical Characterization
This subsection estimates the main moments and the PDF of the MPM parameters, computed for the overall set of measured channels, that have been obtained using the fitting procedure presented in Section III. For each empirical (estimated) PDF, the probability distribution that gives the best fit is given. To this end, the most common distribution functions are considered: beta, Birnbaum-Saunders, exponential, gamma, generalized extreme value (GEV), Gumbel, inverse Gaussian, logistic, log-logistic, lognormal, Nakagami, normal, Poisson, Rayleigh, Rician, $t$ location-scale and Weibull. For each distribution, the parameters that yield the best fit to the set of MPM parameters are obtained by means of the maximum likelihood (ML) estimation assuming independent samples [Bickel2015]. Afterwards, the Anderson-Darling goodness-of-fit test is applied and the distribution yielding the highest $p$-value is selected. When the test gives the same $p$-value for all distributions or the null hypothesis is rejected for all of them, the one with the highest value of the likelihood function is selected. For some parameters, more elaborate models consisting of mixture distributions have been considered.
The mean and standard deviation of the MPM parameters are given in Table IV. The fitting statistical distributions with their corresponding parameter values are summarized in Table V. These reported results allow future replicability of this study and simplify the development of improved statistical models.
Table IV: Mean and standard deviation of the MPM parameters
@@
Figure 4: Empirical and fitted PDFs of the attenuation coefficients $a_{0}$ and $a_{1}$ and the normalization coefficient $A$.
@@
Figure 5: Empirical and fitted PDFs of the number of dominant paths, $N$, the modulus of the path gains $|g_{i}|$ and the path lengths $d_{i}$.
@@
The empirical s of the MPM parameters are reported in Fig. 4 and Fig. 5. Regarding $a_{0}$ and $a_{1}$, whose EPDFs are show in Fig. 4 (a) and (b), respectively, the distribution that gives the best fit is the GEV ($p$-value=0.286=0.286 for $a_{0}$ and 0.8941 for $a_{1}$), although the normal distribution also gives a high $p$-value (0.1967 for $a_{0}$ and 0.5664 for $a_{1}$). In practice, the GEV is a family of three types of extreme value distributions, namely type I, or Gumbel (or log-Weibull) [Gumbel1954], type II, or Fréchet, and type III, or Weibull distribution. This result makes sense since the GEV is the limit distribution of properly normalized maxima of a sequence of independent and identically distributed random variables. The GEV distribution for $a_{0}$ is justified since it represents the maximum (or the minimum in case of negative values of $a_{1}$) of the channel attenuation profile for the identified trend of each measured CFR.
@@
Regarding the normalization coefficient, $A$, whose EPDF is displayed in Fig. 4 (c), the Log-logistic gives the largest $p$-value (0.909) although, although the GEV and the lognormal also give high values ($p$-values $0.788$ and $0.4536$, respectively). The latter is not surprising, as $A$ is determined as the maximum of the absolute value of the path gains, $|g_{i}|$, of each fitted channel.
@@
Regarding the probability distribution of the number of dominant paths, $N$, while this magnitude can only take non-negative integer values, discrete distributions like the binomial and Poisson ones give very poor fitting to the actual set of values. Hence, $N$ has been considered as a discretized version of a continuous random variable. This same approach has been also applied to fit the path length, $d_{i}$. By doing so, the GEV gives the best fit ($p$-value=0.86752=0.86752) to the number of dominant paths, $N$, whose EPDF is displayed in Fig. 5 (a). As a matter of fact, $N$ represents the maximum number of dominant paths that actively contribute to reconstruct each measured CFR.
@@
Concerning the characterization of the path gain, its absolute value, $|g_{i}|$, is considered for the sake of simplicity, as the EPDF of $g_{i}$ shows a symmetric behavior around the $y$-axis. The statistical characterization of $|g_{i}|$ must take into account that the normalization process used to force $g_{i}\in[-1,1]$ in each fitted channel causes the probability of $|g_{i}|=1$ to be non-zero. Hence, $|g_{i}|$ has to be modeled by a mixed random variable (RV) consisting of a continuous component and a discrete term at $|g_{i}|=1$. Fig. 5 (b) shows that the continuous term of the EPDF can be fitted by a lognormal RV. Accordingly, the overall PDF can be expressed as $f(|g_{i}|)=\pi_{0}q(|g_{i}|;\mu,\sigma)+\pi_{1}\delta(|g_{i}|-1)$, where $q(\cdot;\mu,\sigma)$ denotes the PDF of a lognormal RV with parameters $\mu$ and $\sigma$ and $\delta(\cdot)$ is the Dirac delta function. The value of these parameters are given in Table V.
@@
The EPDF of $d_{i}$, displayed in Fig. 5 (c), exhibits a bimodal behavior, with maxima around $d_{i}=50$ m and $d_{i}=3195$ m. The latter is an effect of the use of the discrete frequency domain version of (1), which obliges to set a limit for the largest path length. Since the observed bimodal behavior cannot be obtained with any of the distributions indicated above, we have resorted to mixture models. It has been found that the Weibull distribution gives a good fitting to the values of $d_{i}\leq 1500$ while the GEV, shifted and reflected about $d_{N-1}=\frac{N-1}{N}L$, adequately fits the values of $d_{i}>1500$. Hence, the PDF of $d_{i}$ can be modeled as $f(d_{i})=\pi_{0}g_{0}(d_{i};\lambda_{0},k_{0})+\pi_{1}g_{1}(d_{N-1}-d_{i};k_{1},\sigma_{1},\mu_{1})$, where $g_{0}(\cdot;\lambda_{0},k_{0})$ denotes the PDF of a Weibull distribution with scale and shape parameters $\lambda_{0}$ and $k_{0}$, respectively, and $g_{1}(\cdot;k_{1},\sigma_{1},\mu_{1})$ is the PDF of a GEV distribution with location, scale and shape parameters $k_{1}$, $\sigma_{1}$ and $\mu_{1}$, respectively, whose values are given in Table V. As shown in Fig. 5 (c), the PDF of the proposed mixture gives a good fitting to the empirical one.
Table V: Fitting distributions of the MPM parameters
@@
In order to get an in-depth analysis of the behavior of $d_{i}$, Fig. 6 (a) displays the path length, $d_{i}$, vs the dominant path index, $i$, for some illustrative measured channels. For the sake of clarity, the markers corresponding the same channel have been linked by a dashed line of the the same color. As seen, all channels have a set of paths whose length is uniformly distributed from zero (or almost zero) and a given value that depends on the channel. However, some channels, like Ch. 3-7, also have a set of paths with very large length. In many cases, like Ch. 3 and Ch. 4, there are no paths with intermediate distances. This behavior is related to the frequency selectivity of the channel: the larger the frequency selectivity, the larger the distances of the paths required to synthesize is. This can be observed in Fig. 6 (b), where the scatter plot of the median value of the dominant path lengths vs the delay spread of each measured channel is given. As observed, channels with higher delay spread values (larger frequency selectivity) need larger paths to be synthesized. Furthermore, it is also expected that the higher the number of dominant paths, the larger the median length of the paths. Fig. 6 (c) confirms this end by showing a strong linear relationship between them.